OpenStudy (anonymous):

(sin x)(tan x cos x - cot x cos x) = 1 - 2 cos2x

Nnesha (nnesha):

rewrite tan and cot in terms of sin cos

OpenStudy (anonymous):

(sin x)(tan x cos x - cot x cos x) = 1 - 2 cos^2x

OpenStudy (anonymous):

(sinx)(sinx/cosx cosx- cosx/sinx cosx)

OpenStudy (anonymous):

what do i do next

Nnesha (nnesha):

\[\large\rm \sin(x)(\frac{\sin x}{cosx} * cosx -\frac{\cos x }{sinx } *\cos x) \] first solve the parentheses sinx/cos times cos x = ?

OpenStudy (anonymous):

sinx

OpenStudy (anonymous):

\[sinx(sinx-\cos^2x /sinx)\]

Nnesha (nnesha):

right cos cancel out \[\large\rm \sin(x)(\frac{\sin x}{\cancel{cosx}} *\cancel{ cosx} -\frac{\cos x }{sinx } *\cos x) \] \[\large\rm \sin(x)( sin x -\frac{\cos x }{sin x } *\cos x) \] now you can distribute by sinx

Nnesha (nnesha):

right

OpenStudy (anonymous):

\[\sin^2x-\cos^2x\]

Nnesha (nnesha):

\[\rm sinx ( sinx - \frac{\cos^2x}{sinx})\] multply both terms in the parentheses by sin x

Nnesha (nnesha):

oh wow nice that's right

OpenStudy (anonymous):

what do i do from there

Nnesha (nnesha):

now we can use the identity \[\huge\rm sin^2 \theta + \cos^2 \theta = 1\] solve this identity for sin^2x

OpenStudy (anonymous):

\[\sin^2x=1-\cos^2x\]

OpenStudy (anonymous):

i get it now thank you

Nnesha (nnesha):

nice !! we can replace sin^2x with 1-cos^2x \[\huge\rm \color{Red}{sin^2x}-\cos^2x\] \[\huge\rm \color{Red}{(1- \cos^2x)}-\cos^2x\] now combine like terms

Nnesha (nnesha):

nice good work ! keep it up! :=))